# Reevaluating Policy Gradient Methods for Imperfect-Information Games

### v2

@@ Abstract
In the past decade, motivated by the putative failure of naive self-play deep reinforcement learning (DRL) in adversarial imperfect-information games, researchers have developed numerous DRL algorithms based on fictitious play (FP), double oracle (DO), and counterfactual regret minimization (CFR). In light of recent results of the magnetic mirror descent algorithm, we hypothesize that simpler generic policy gradient methods like PPO are competitive with or superior to these FP-, DO-, and CFR-based DRL approaches. To facilitate the resolution of this hypothesis, we implement and release the first broadly accessible exact exploitability computations for four large games. Using these games, we conduct the largest-ever exploitability comparison of DRL algorithms for imperfect-information games. Over 5600 training runs, we find that FP-, DO-, and CFR-based approaches fail to outperform generic policy gradient methods.
Code available at https://github.com/nathanlct/IIG-RL-Benchmark and https://github.com/gabrfarina/exp-a-spiel.
@@ 1 Introduction
Unfortunately, constructing effective DRL approaches from these tabular algorithms has proven challenging. FP and DO require an expensive best response computation at each iteration (i.e., each iteration requires solving an entire reinforcement learning problem) and can exhibit slow convergence as a function of iteration number (Zhang and Sandholm, 2024). While CFR converges more quickly, its immediate adaptation to model-free reinforcement learning requires importance weighting over trajectories (Lanctot et al., 2009), resulting in high variance feedback to which function approximation is poorly suited. Moreover, none of FP, DO and CFR enjoy last-iterate convergence, necessitating a layer of indirection to extract the desired policy.
@@
However, assessing the above hypothesis is not straightforward. Doing so demands exact exploitability computations for games large enough to be representative of settings to which DRL is relevant, but small enough that these computations are tractable in reasonable time. Unfortunately, due to the complexity of implementing these computations for large games, few are publicly available, even in OpenSpiel (Lanctot et al., 2019)—the standard library for DRL in IIGs. This issue has plagued the field, forcing researchers to resort to tenuous metrics and making progress difficult to measure.
We address this issue by implementing exact exploitability computation for four large games: 3x3 Dark Hex, Phantom Tic-Tac-Toe, 3x3 Abrupt Dark Hex, and Abrupt Phantom Tic-Tac-Toe. These games have 10s of billions of nodes in their game trees—over one million times more than many games for which exploitability is typically reported, such as Leduc Hold’em (Southey et al., 2005).
Using these benchmarks, we undertake the largest-ever exploitability comparison of DRL algorithms for IIGs, spanning 5600 total runs across 350 hyperparameter configurations, requiring over 278,000 CPU hours, and including NFSP (Heinrich and Silver, 2016), PSRO (Lanctot et al., 2017), ESCHER (McAleer et al., 2023), R-NaD (Perolat et al., 2022), MMD (Sokota et al., 2023), PPO (Schulman et al., 2017) and PPG (Cobbe et al., 2021). Over these runs, all of which are included in the paper, NFSP, PSRO, ESCHER, and R-NaD fail to outperform the generic PG methods (MMD, PPO, PPG).
@@ 2.1 Game Formalism
IIGs are often formalized as perfect-recall extensive-form games (EFGs) (Kuhn, 2003). Our formalism is equivalent to perfect-recall EFGs, but possesses greater superficial similarity to traditional reinforcement learning notation. We notate a game as a tuple
$\langle\mathbb{S},\mathbb{A},\mathbb{O},\mathbb{I},\mathcal{R},\mathcal{T},\mathcal{O},\mathcal{C},t_{\text{max}}\rangle,$
where
  • $\mathbb{S}$ is the space of game states;
  • $\mathbb{A}$ is the space of actions;
  • $\mathbb{O}$ is the space of observations;
  • $\mathbb{I}=\cup_{t}(\mathbb{O}\times\mathbb{A})^{t}\times\mathbb{O}$ is the space of information states;
  • $\mathcal{R}\colon\mathbb{S}\times\mathbb{A}\to\mathbb{R}$ is the reward function for player 1 (and the cost function for player 2);
  • $\mathcal{T}\colon\mathbb{S}\times\mathbb{A}\to\Delta(\mathbb{S})$ is the transition function;
  • $\mathcal{O}\colon\mathbb{S}\to\mathbb{O}$ is the observation function, which determines the observation given to the acting player;
  • $\mathcal{C}\colon\mathbb{S}\to\{1,2\}$ is the player choice function, which determines the acting player;
  • $t_{\text{max}}$ is the time step at which the game terminates.
@@
The objective of player 1 and loss of player 2 is the expected return
$\mathcal{J}(\pi)=\mathbb{E}_{\pi}\left[\sum_{t=1}^{t_{\text{max}}}\mathcal{R}(S^{t},A^{t})\right].$
However, directly optimizing the expected return is only possible given a particular opponent or distribution of opponents. In practice, identifying this distribution may be prohibitively expensive or altogether infeasible—such a distribution may not even exist.
@@ 2.2.1 Naive self-play
Of these approaches, naive self-play is appealing for its simplicity; however, it is easy to show that it can produce catastrophic results. Some value-based algorithms cannot even express non-deterministic policies. This makes it impossible for them to achieve low-exploitability in many games (e.g., in rock-paper-scissors, deterministic policies are maximally exploitable). While PG methods can at least express non-deterministic policies, their learning dynamics generally cycle, diverge or exhibit chaotic behavior, rather than converge to Nash equilibria (Cheung and Piliouras, 2019).
@@ 2.2.2 Best Responses
An alternative is to leverage tabular game-theoretic algorithms that rely on best-response computations. These game-theoretic algorithms support a plug and play workflow with model-free DRL, by substituting the exact best response for an approximate best response computed via DRL. And, unlike naive self-play, they come with convergence guarantees.
@@
Unfortunately, both FP and DO can suffer from slow convergence. This is suggested not only by practice, but also according to the exponential best-known complexity for FP and exponential lower bound for DO (Zhang and Sandholm, 2024). Requiring a large number of iterations is especially concerning in the context of DRL because each iteration requires solving an entire reinforcement learning problem. Despite numerous works on DRL variants of both FP (Heinrich and Silver, 2016; Qin et al., 2019; Vinyals et al., 2019; Goldstein and Brown, 2023; Hu et al., 2023; Li et al., 2023a) and DO (Lanctot et al., 2017; McAleer et al., 2020, 2021; Smith et al., 2021; McAleer et al., 2022; Dinh et al., 2022; Zhou et al., 2022; Yao et al., 2023; Tang et al., 2023; McAleer et al., 2024; Li et al., 2024a; Lian, 2024), it is not clear that this issue has been addressed.
@@ 2.2.3 Counterfactual Regret Minimization
Because CFR is more involved than FP and DO, it does not admit the same kind of plug and play workflow with single-agent model-free DRL. Instead, new methods (Steinberger et al., 2020; Gruslys et al., 2020; McAleer et al., 2023) have been designed specifically around CFR. Unfortunately, these new methods come with new weaknesses. The most straightforward method (Steinberger et al., 2020), for instance, relies on importance sampling over trajectories, which leads to high variance in feedback, especially in games with many steps. This makes it challenging for function approximators to learn. Although McAleer et al. (2023) resolve this issue, their resolution necessitates a uniform behavioral policy, making it difficult to achieve sufficient coverage of relevant trajectories.
@@ 2.2.4 Regularization
[…]
@@ 3 The Policy Gradient Hypothesis
  • $-\alpha\text{KL}(\pi,\rho)$ regularizes the updated policy (for uniform $\rho$, this acts similarly to an entropy bonus).
  • $-\frac{1}{\eta}\text{KL}(\pi,\pi_{t})$ constrains the update size.
@@
A second possibility is that these generic PG methods are fully capable of performing well, but have not been run with good hyperparameters. Ever-present reasons that baselines tend to perform poorly, such as the expense and difficulty of tuning (and, not least, the disinclination of researchers toward showing baselines outperforming their algorithmic contributions), could be at work.
@@
The confirmation of the policy gradient hypothesis would have ramifications on both research and practice. On the research side, it would vindicate generic PG methods, which have generally either been dismissed as unsound or relegated to the role of sacrificial baseline. Practically, it would lead to simpler implementations and potentially significant performance improvements in applications where FP-, DO-, and CFR-based DRL algorithms are currently employed. These applications include autonomous vehicles (Ma et al., 2018), network security (Lu et al., 2020; Xue et al., 2021), robot confrontation (Hu et al., 2022), eavesdropping (Guo et al., 2022), radar (anti-)jamming (Li et al., 2022a, b; Geng et al., 2023, 2024), intrusion response (Hammar and Stadler, 2023), aerial combat (He et al., 2023; Wang et al., 2024), pursuit-evasion games (Li et al., 2023b, 2024b), racing (Zheng et al., 2024), language model alignment (Gemp et al., 2024) and red teaming (Ma et al., 2024).
@@ 4 Benchmarks
Unfortunately, evaluating DRL algorithms in adversarial IIGs is not straightforward because of the dearth of efficient exploitability computations for large IIGs. As a result of this dearth, only a small number of works in the literature (Steinberger et al., 2020; Gruslys et al., 2020; Hu et al., 2023) make the effort to report exploitability in large games.
The rest of the literature predominantly resorts to some combination of two unsatisfactory metrics. The first is exploitability in the same small EFGs that served as benchmarks for tabular solvers, such as Leduc Hold’em (Southey et al., 2005). Benchmarking in these small games is unsatisfactory because success requires learning a precise near-Nash policy for a handful of repeatedly revisited information states—a fundamentally different challenge from large games, where success requires learning a policy that is strong (but not necessarily numerically close to Nash) for a vast number of unvisited information states.
The second is head-to-head evaluations in large games. These evaluations are unsatisfactory because the intransitive dynamics typically present in imperfect-information games, such as rock-beats-scissors-beats-paper-beats-rock, muddle interpretations of relative strength. Exacerbating the issue, there is no community coordination on head-to-head opponents, precluding direct cross-paper comparisons.
This predicament has created an undesirable state of affairs. Many works claim “state-of-the-art” performance, but even experts in specific algorithm classes are often unsure which instances are actually effective.
To address this predicament, we set out to implement efficient computations for four large games—Phantom Tic-Tac-Toe (PTTT), 3x3 Dark Hex (DH3), Abrupt Phantom Tic-Tac-Toe (APTTT), and 3x3 Abrupt Dark Hex (ADH3)—ourselves. These games are imperfect-information twists on the perfect-information alternating-moves games Tic-Tac-Toe and 3x3 Hex.
Figure 1: A game of Tic-Tac-Toe (top) and 3x3 Hex (bottom).
@@
    OpenSpiel (Lanctot et al., 2019) support: DH3, PTTT, and ADH3 are already implemented in OpenSpiel, making it easy for other researchers to use them. To address APTTT’s absence, we provide our own game implementation.
@@
    Precedent: Existing work uses PTTT and ADH3 for head-to-head evaluations (Sokota et al., 2023; McAleer et al., 2023; Meng et al., 2023; Liu et al., 2024), providing external evidence of their relevance.
@@
    Size: These games allow model-free DRL algorithms to be trained in minutes or hours on commodity hardware. But, as detailed in Table 1, they have 10s of billions of game states, making them quite large by the standards of game solving benchmarks.
Table 1: Fundamental game quantities. Positive Nash values (i.e., expected values for player 1 at Nash equilibria) mean that player 1 possesses a structural advantage. The Nash value of 1 for DH3 implies a guaranteed winning strategy for player 1, which we describe in Appendix B.
Game States Info. states Nash value $\pm$ error
DH3 19.12B 6.07M 1.$0$$\pm$ 0.00000
ADH3 29.31B 27.33M 0.38443 $\pm$ 0.00021
PTTT 19.93B 5.99M 0.66665 $\pm$ 0.00026
APTTT 27.12B 23.31M 0.55017 $\pm$ 0.00014
Indeed, the size of these games makes it nontrivial to implement exploitability computations. To accommodate their magnitude, we utilize sequence-form representations, which are approximately 1000 times more compact than the explicit game trees, alongside dynamic payoff matrix generation to circumvent explicit matrix storage. This memory-efficient approach enables usage on commodity hardware.
To optimize runtime, we distribute computation across multiple threads after the first two moves, using 18 buffers to prevent concurrency conflicts across the 81 possible opening sequences. Depending on the machine and game, this implementation delivers exploitability computation (and head-to-head evaluation) times of roughly 30 to 90 seconds.

### v3

@@ Abstract
In the past decade, motivated by the putative failure of naive self-play deep reinforcement learning (DRL) in adversarial imperfect-information games, researchers have developed numerous DRL algorithms based on fictitious play (FP), double oracle (DO), and counterfactual regret minimization (CFR). In light of recent results of the magnetic mirror descent algorithm, we hypothesize that simpler generic policy gradient methods like PPO are competitive with or superior to these FP-, DO-, and CFR-based DRL approaches. To facilitate the resolution of this hypothesis, we implement and release the first broadly accessible exact exploitability computations for five large games. Using these games, we conduct the largest-ever exploitability comparison of DRL algorithms for imperfect-information games. Over 7000 training runs, we find that FP-, DO-, and CFR-based approaches fail to outperform generic policy gradient methods.
footnotetext: Code available at https://github.com/nathanlct/IIG-RL-Benchmark and https://github.com/gabrfarina/exp-a-spiel.
@@ 1 Introduction
Unfortunately, constructing effective DRL approaches from these tabular algorithms has proven challenging. FP and DO require expensive best response computations at each iteration (i.e., each iteration requires solving an entire reinforcement learning problem) and can exhibit slow convergence as a function of iteration number (Zhang and Sandholm, 2024). While CFR converges more quickly, its immediate adaptation to model-free reinforcement learning requires importance weighting over trajectories (Lanctot et al., 2009), resulting in high variance feedback to which function approximation is poorly suited. Moreover, none of FP, DO and CFR enjoy last-iterate convergence, creating a layer of indirection to extract the desired policy.
@@
However, assessing the above hypothesis is not straightforward. Doing so demands exact exploitability computations for games large enough to be representative of settings to which DRL is relevant, but small enough that such computations are efficient. Unfortunately, due to the complexity of implementing these computations for large games, few are publicly available, even in OpenSpiel (Lanctot et al., 2019)—the standard library for DRL in IIGs. This issue has plagued the field, forcing researchers to resort to tenuous metrics and making progress difficult to measure.
We address this issue by implementing exact exploitability computation for five large games: 2D5F Liar’s Dice, 3x3 Dark Hex, Phantom Tic-Tac-Toe, 3x3 Abrupt Dark Hex, and Abrupt Phantom Tic-Tac-Toe. These games have millions of infostates—tens of thousands more than many games for which exploitability is typically reported, such as Leduc Hold’em (Southey et al., 2005).
Using these benchmarks, we undertake the largest-ever exploitability comparison of DRL algorithms for IIGs, spanning 7000 total runs across 350 hyperparameter configurations, requiring over 345,000 CPU hours, and including NFSP (Heinrich and Silver, 2016), PSRO (Lanctot et al., 2017), ESCHER (McAleer et al., 2023), R-NaD (Perolat et al., 2022), MMD (Sokota et al., 2023), PPO (Schulman et al., 2017) and PPG (Cobbe et al., 2021). Over these runs, all of which are included in the paper, NFSP, PSRO, ESCHER, and R-NaD fail to outperform the generic PG methods (MMD, PPO, PPG).
@@ 2.1 Game Formalism
IIGs are often formalized as perfect-recall extensive-form games (EFGs) (Kuhn, 2003). Our formalism is equivalent to perfect-recall EFGs, but possesses greater superficial similarity to traditional reinforcement learning notation. We notate a game as a tuple $\langle\mathbb{S},\mathbb{A},\mathbb{O},\mathbb{I},\mathcal{R},\mathcal{T},\mathcal{O},\mathcal{C},t_{\text{max}}\rangle$, where $\mathbb{S}$ is the set of game states; $\mathbb{A}$ is the set of actions; $\mathbb{O}$ is the set of observations; $\mathbb{I}=\cup_{t}(\mathbb{O}\times\mathbb{A})^{t}\times\mathbb{O}$ is the set of information states; $\mathcal{R}\colon\mathbb{S}\times\mathbb{A}\to\mathbb{R}$ is the reward function for player 1 (and the cost function for player 2); $\mathcal{T}\colon\mathbb{S}\times\mathbb{A}\to\Delta(\mathbb{S})$ is the transition function; $\mathcal{O}\colon\mathbb{S}\to\mathbb{O}$ is the observation function, which determines the acting player’s observation; $\mathcal{C}\colon\mathbb{S}\to\{1,2\}$ is the player choice function, which determines the acting player; $t_{\text{max}}$ is the time step at which the game terminates.
@@
The objective of player 1 and loss of player 2 is the expected return $\mathcal{J}(\pi)=\mathbb{E}_{\pi}\sum\nolimits_{t=1}^{t_{\text{max}}}\mathcal{R}(S^{t},A^{t})$.
Directly optimizing the expected return is only possible given a particular opponent or distribution of opponents. In practice, identifying this distribution may be prohibitively expensive or altogether infeasible—such a distribution may not even exist.
@@ 2.2.1 Naive self-play
Of these approaches, naive self-play is appealing for its simplicity; however, it is easy to show that it can produce catastrophic results. Some value-based algorithms cannot even express non-deterministic policies. This makes it impossible for them to achieve low-exploitability in many games (e.g., in rock-paper-scissors, deterministic policies are maximally exploitable). While PG methods can express non-deterministic policies, their learning dynamics generally cycle, diverge or exhibit chaotic behavior, rather than converge to Nash equilibria (Cheung and Piliouras, 2019).
@@ 2.2.2 Best Responses
A next-simplest alternative is to leverage tabular game-theoretic algorithms that rely on best-response computations. These game-theoretic algorithms support a plug and play workflow with model-free DRL, by substituting the exact best response for an approximate best response computed via DRL. And, unlike naive self-play, they come with convergence guarantees.
@@
Unfortunately, both FP and DO can suffer from slow convergence. Requiring a large number of iterations is especially concerning in the context of DRL because each iteration requires solving an entire reinforcement learning problem. Despite numerous works on DRL variants of both FP (Heinrich and Silver, 2016; Qin et al., 2019; Vinyals et al., 2019; Goldstein and Brown, 2023; Hu et al., 2023; Li et al., 2023a) and DO (Lanctot et al., 2017; McAleer et al., 2020; 2021; Smith et al., 2021; McAleer et al., 2022b; Dinh et al., 2022; Zhou et al., 2022; Yao et al., 2023; Tang et al., 2023; McAleer et al., 2024; Li et al., 2024b; Lian, 2024), it is not clear that this issue has been addressed.
@@ 2.2.3 Counterfactual Regret Minimization
Because CFR is more involved than FP and DO, it does not admit the same plug and play workflow with single-agent model-free DRL. Instead, new methods (Steinberger et al., 2020; Gruslys et al., 2020; McAleer et al., 2023) were designed specifically for CFR. Unfortunately, these new methods come with new weaknesses. The most straightforward method (Steinberger et al., 2020), for instance, relies on importance sampling over trajectories, which leads to high variance in feedback, especially in games with many steps. This makes it challenging for function approximators to learn. Although McAleer et al. (2023) resolve this issue, their resolution necessitates a uniform behavioral policy, making it difficult to achieve sufficient coverage of relevant trajectories.
@@ 2.2.4 Regularization
$\pi_{t+1}=\underset{\pi}{\text{argmax}}\underset{A\sim\pi}{\mathbb{E}}q(A)-\alpha\text{KL}(\pi,\rho)-\frac{1}{\eta}\text{KL}(\pi,\pi_{t}),$ (1)
@@ 3 The Policy Gradient Hypothesis
  • $\alpha\text{KL}(\pi,\rho)$ regularizes the updated policy (for uniform $\rho$, this acts similarly to an entropy bonus).
  • $\frac{1}{\eta}\text{KL}(\pi,\pi_{t})$ constrains the update size.
@@
A second possibility is that these generic PG methods are fully capable of performing well, but have not been run with good hyperparameters. Ever-present reasons that baselines tend to perform poorly, such as the expense and difficulty of tuning and the structural disincentives to reporting strong results for baselines, could be at work.
@@
The confirmation of the policy gradient hypothesis would have ramifications on both research and practice. On the research side, it would vindicate generic PG methods, which have generally either been dismissed as unsound or relegated to the role of sacrificial baseline. Practically, it would lead to simpler implementations and performance improvements in applications where FP-, DO-, and CFR-based DRL algorithms are currently employed. These applications include autonomous vehicles (Ma et al., 2018), network security (Lu et al., 2020; Xue et al., 2021), robot confrontation (Hu et al., 2022), eavesdropping (Guo et al., 2022), radar (anti-)jamming (Li et al., 2022b; a; Geng et al., 2023; 2024), intrusion response (Hammar and Stadler, 2023), aerial combat (He et al., 2023; Wang et al., 2024), pursuit-evasion games (Li et al., 2023b; 2024a), racing (Zheng et al., 2024), language model alignment (Gemp et al., 2024) and red teaming (Ma et al., 2024).
@@ 4 Benchmarks
Unfortunately, evaluating DRL algorithms for adversarial IIGs is not well standardized. Due to the difficulty of efficiently implementing exploitability for large IIGs, there is a dearth of accessible implementations.
As a result, the literature predominantly resorts to some combination of two unsatisfactory metrics. The first is exploitability in the same small EFGs that served as benchmarks for tabular solvers, such as Leduc Hold’em (Southey et al., 2005) (see Table 4). Benchmarking in these small games is unsatisfactory because success requires learning a precise near-Nash policy for a handful of repeatedly revisited information states—a fundamentally different challenge from large games, where success requires learning a policy that is strong (but not necessarily numerically close to Nash) for a vast number of unvisited information states.
The second is head-to-head evaluations in large games. These evaluations are unsatisfactory because the intransitive dynamics characteristic of imperfect-information games (e.g., rock-beats-scissors-beats-paper-beats-rock), muddle interpretations of relative strength. Exacerbating the issue, there is no community coordination on head-to-head opponents, precluding direct cross-work comparisons.
This predicament has created an undesirable state of affairs. Many works claim “state-of-the-art” performance, but even experts in specific algorithm classes are often unsure which instances are actually effective. To address this predicament, we set out to implement efficient computations for five large games: 2D5F Liar’s Dice (LD2D5F), Phantom Tic-Tac-Toe (PTTT), 3x3 Dark Hex (DH3), Abrupt Phantom Tic-Tac-Toe (APTTT), 3x3 Abrupt Dark Hex (ADH3).
In 2D5F Liar’s Dice, each of two players privately rolls two dice with five faces. Players then alternate making claims about the total number of dice across both players that show a given face value (e.g., “two dice showing four”). A legal bid (in the ruleset OpenSpiel adopts) must either increase the quantity or increase the face value while not decreasing the quantity. On a turn, instead of bidding, a player may also call “liar,” whereupon all dice are revealed; if the last bid is no greater than the revealed count of the specified face, the the bidder wins, otherwise the caller does.
Figure 1: Tic-Tac-Toe (left) and 3x3 Hex (right).
Table 1: Game quantities. Positive Nash values (i.e., expected values for player 1 at Nash equilibria) mean that player 1 has a structural advantage.
Game States Info. states Nash value
LD2D5F 235.93M 15.73M $0.01770\pm 1.8e^{-4}$
DH3 19.12B 6.07M $1.00000\pm 0.0$$e^{-4}$
ADH3 29.31B 27.33M $0.38443\pm 2.1e^{-4}$
PTTT 19.93B 5.99M $0.66665\pm 2.6e^{-4}$
APTTT 27.12B 23.31M $0.55017\pm 1.4e^{-4}$
@@
    OpenSpiel (Lanctot et al., 2019) support: LD2D5F, DH3, PTTT, ADH3 are already implemented in OpenSpiel, making it easy for other researchers to use them. To address APTTT’s absence, we provide our own game implementation.
@@
    Precedent: Existing work uses LD, PTTT, and ADH3 for head-to-head evaluations (Sokota et al., 2023; McAleer et al., 2023; Meng et al., 2023; Liu et al., 2024), providing external evidence of their relevance.
@@
    Size: These games allow model-free DRL algorithms to be trained in minutes or hours on commodity hardware. But, as detailed in Table 1, they have millions of infostates, making them quite large by the standards of game solving benchmarks.
Indeed, the size of these games makes it nontrivial to implement exploitability computations. To accommodate their magnitude, we use sequence-form representations, which are approximately multiple orders of magnitude more compact than the explicit game trees, alongside dynamic payoff matrix generation to circumvent explicit matrix storage. This memory-efficient approach enables usage on commodity hardware.
To optimize the runtime of LD2D5F, we exploit the facts that actions are public and chance outcomes are independent and located at the root of the game to use dynamic programming to calculate the payoff at all public states in linear time, similarly to Brown et al. (2020). To optimize the runtime of the 3x3 games, we distribute computation across multiple threads after the first two moves, using 18 buffers to prevent concurrency conflicts across the 81 possible opening sequences. Depending on the machine and game, this implementation delivers exploitability computation (and head-to-head evaluation) times of roughly 30 to 90 seconds.